OpenStudy (anonymous):

2 +3 = 3+ 2 On the lines below, explain how the completed equation demonstrates the commutative property of addition. 2 +3 = 3+ 2 On the lines below, explain how the completed equation demonstrates the commutative property of addition. @Mathematics

OpenStudy (anonymous):

it is not a "property", it is an axiom. Axiom is a proposition accepted as true, without demostration.

OpenStudy (anonymous):

2+3 = 3+2 5 = 5 As you can see, in whatever order you add the two numbers, you get 5. So the equation demonstrates the commutative property of addition.

OpenStudy (anonymous):

it is not a "property", it is an axiom. Axiom is a proposition accepted as true, without demostration